simplify/solve it reads x-1 over x-3 is equal to 2 over x-3:

x-1 2
--- = ---
x-3 x-3

Answer by checkley77(12844)

x-1 2
--- = --- cross multiply
x-3 x-3
(x-1)(x-3)=2(x-3)
x^2-4x+3=2x-6
x^2-4x-2x+3+6=0
x^2-6x+9=0
(x-3)(x-3)=0
x-3=0
x=3 answer.
proof
(3-1)/3-3)=2/(3-3)
2/0=2/0
0=0